OpenStudy (anonymous):

CJ will use 3 bags of cement for his project. Each bag of cement needs 3,1/2 gallons of water added to it. How much water will CJ need? Express your answer in simplest form. A. 9, 1/2 gallons B. 10 ,1/2 gallons C. 11 gallons D. 11,1/2 gallons

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats for all of them

OpenStudy (anonymous):

"FAN AND MEDAL"

OpenStudy (anonymous):

\[\frac{3\ 7}{2}=\frac{21}{2}=10+\frac{1}{2} \] The input for 3 space 7 / 2, was 3 times 7/2

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

@hhelpplzzzz

OpenStudy (anonymous):

@IHelpYouLearn

OpenStudy (anonymous):

i have 9 q more